# Scalarisation-based risk concepts for robust multi-objective optimisation

### v2

@@ 7 Conclusion
In this work, we studied the robust multi-objective optimisation problem from a computational perspective. We identified that the majority of existing robust multi-objective optimisation algorithms, in one way or another, rely on two key operations: robustification and scalarisation. These two procedures are in general not commutative and therefore the order that they are applied in has an effect on the resulting solutions that are identified and the final decisions that are made. As part of our work, we analysed the effect of this ordering from both a philosophical and a methodological perspective. In particular, after setting up the problem in Sections 2 and 3, we presented a useful discussion on the different scenarios when one should prefer one ordering over another (Section 3.3.2) and when these approaches are equivalent (Section 3.3.3). We then moved on to the next sections where we described how these two different orderings can be used in order to extend classical concepts and tools from standard multi-objective optimisation to the robust setting. Specifically, we defined the general concept of a robust Pareto front (Section 4) and a robust performance metric (Section 5). We then demonstrated the utility of these new ideas via some insightful numerical experiments in Section 6.
@@ Appendix A Proofs
due to the weak optimality of $\mathbf{y}^{*}$ and the strong monotonicity of the length scalarisation function (Tu et al., 2024b, Lemma A.2). This further implies that
@@
The proof of this result is presented by Tu et al. (2024b, Appendix A.1).
@@ A.5 Pareto front conditions
In order to verify whether a polar surface is a Pareto front surface, we will rely on the key result by Tu et al. (2024b, Proposition 3.1), which we recall below in Proposition A.1.
@@
(Tu et al., 2024b, Proposition 3.1) Consider a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, then this set is a Pareto front surface, that is $A=\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]$, if and only if the following two conditions holds:
@@
Now consider the setting where there exists one direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$ in which the projected length (28) is zero. This implies that $\max_{\mathbf{x}\in\mathbb{X}}s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(f(\mathbf{x},\boldsymbol{\xi}))$ is zero on the support of $\rho$. As the univariate risk functional is partially coherent and the length scalarisation function is strongly monotonic (Tu et al., 2024b, Lemma A.2), this implies that all other directions must also have a projected length of zero. Therefore, in this setting, the polar surface (19) reduces down to the degenerate singleton set.
@@ A.10 Hypervolume of a Pareto front surface
The proof of Proposition 5.1 follows directly from the definition of the RTS front (20), together with Proposition 4.2, Lemma A.1 and the monotonicity of the hypervolume transformation. The proof of Proposition 5.2 follows in a similar way.

### v4

@@ 7 Conclusion
In this work, we studied the robust multi-objective optimisation problem from a computational perspective. We identified that the majority of existing robust multi-objective optimisation algorithms, in one way or another, rely on two key operations: robustification and scalarisation. These two procedures are in general not commutative and therefore the order that they are applied in has an effect on the resulting solutions that are identified and the final decisions that are made. As part of our work, we analysed the effect of this ordering from both a philosophical and a methodological perspective. Notably, among other things, we proposed the novel concepts of a robust Pareto front surface and a robust performance metric. We then demonstrated the practical utility of these new ideas via some insightful numerical experiments.
@@ Appendix A Proofs
due to the weak optimality of $\mathbf{y}^{*}$ and the strong monotonicity of the length scalarisation function (Tu et al., 2024, Lemma A.2). This further implies that
@@
The proof of this result is presented by Tu et al. (2024, Appendix A.1).
@@ A.5 Pareto front conditions
In order to verify whether a polar surface is a Pareto front surface, we will rely on the key result by Tu et al. (2024, Proposition 3.1), which we recall below in Proposition A.1.
@@
(Tu et al., 2024, Proposition 3.1) Consider a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, then this set is a Pareto front surface, that is $A=\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]$, if and only if the following two conditions holds:
@@
Now consider the setting where there exists one direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$ in which the projected length (30) is zero. This implies that $\max_{\mathbf{x}\in\mathbb{X}}s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(f(\mathbf{x},\boldsymbol{\xi}))$ is zero on the support of $\rho$. As the univariate risk functional is partially coherent and the length scalarisation function is strongly monotonic (Tu et al., 2024, Lemma A.2), this implies that all other directions must also have a projected length of zero. Therefore, in this setting, the polar surface (21) reduces down to the degenerate singleton set.
@@ A.10 Hypervolume of a Pareto front surface
The proof of Proposition 5.1 (or Proposition 5.2) follows directly from the definition of the RTS front (22) (or STR front (23)), together with Proposition 4.2 (or Proposition 4.3), Lemma A.1 and the monotonicity of the hypervolume transformation.